OpenStudy (anonymous):

how do you solve e^-t/10=1/10

OpenStudy (anonymous):

take ln of both sides --> -t/10 = ln(1/10) --> t = -10ln(0.1)

OpenStudy (anonymous):

okay, thanks. how do i do cos 5pi/4? or costheta=1/2

OpenStudy (anonymous):

how do i differentiate y=sin (2x+pi/4) when x=0 to find the tangent and the normal?

OpenStudy (anonymous):

also, how do i find the equation of the normal to y=ln(x+2) which is parallel to the line with the equation, y+3x-5=0, i know it has to have the same gradient which is -3 but i don't know how to go from there.